write a quadratic equation with the given solutions. Type the equation in standard form. ax^2+bx+c=0
the soultions are -11,1.

Answer by Theo(13342)

the solutions are x = -11 or x = 1
this gets you factors of (x+11) * (x-1) = 0
multiply these factors out to get:
x^2 - x + 11x - 11 = 0
combine like terms to get:
x^2 + 10x - 11 = 0
that's your equation in standard form.